# What To Do (and Not to Do) with Causal Panel Analysis under Parallel Trends:

### v2

@@ 5.  Systematic Assessment
The encouraging news is that in over 40% of the studies, we have fairly substantial evidence to attest that the PTA is plausible. In these instances, the DTE estimates in the pre-treatment periods align with zero, and formal statistical tests suggest that the remaining imbalance likely results from randomness in data. Specifically, in these papers the $F$ test does not reject and the equivalence test does reject using the $|\widehat{ATT}|$ as the equivalence threshold, both at the 5% level. In approximately 20% of studies, we find compelling evidence that the PTA is implausible. These are the cases in which we observe a strong pretrend and statistical tests indicate violations of the PTA, i.e. the $F$ test rejects at the 5% level. For the remaining 40% of studies, the results are less clear: We either lack the statistical power or a sufficient number of pretreatment periods to reliably assess the plausibility of the PTA.
HTE-robust estimators do not alter substantive findings but affect power.
To examine the possible impact of the weighting problem caused by HTE, we compared the estimates obtained from the imputation estimator, FEct, to those originally reported. The dissimilarity between the FEct estimates and the original estimates proxies the extent to which the negative weighting problem is consequential in practice.
Figure 5 plots the comparison. The horizontal axis represents the originally reported TWFE estimates, and the vertical axis represents estimates from FEct, both normalized using the originally reported SE. If the point estimates are identical, then the corresponding point should lie exactly on the 45-degree line (dashed). Red solid circles represent studies whose FEct estimates are not significant at the 5% level based on bootstrapped CIs. To assess how the imputation estimator differs from TWFE, we depict the ratio of the FEct estimates to the TWFE estimates, after excluding the always-treated units (ensuring identical sample sets), using solid circles in Figure 6. We also juxtapose the cited TWFE estimates with those omitting the always-treated units, represented by hollow circles. The red circles denote studies where the FEct estimates are not statistically significant at the 5% level.
Figure 5: TWFE vs. Imputation Method
Note: The above figures compare TWFE coefficients and estimates from the imputation method (FEct). Estimates for each application are normalized by the same TWFE SE. The red solid circles represent studies whose ATT estimates from FEct are statistically insignificant at the 5% level based on bootstrapped CIs. Fouirnaies and Hall (2018) and Hall and Yoder (2022) are close to the 45-degree line but not included in the figure as their TWFE $z$-scores are above 15.
We make several observations on Figures 5 and 6. First and foremost, the FEct estimates and the original estimates are highly correlated and always bear the same signs. This suggests that scenarios where accounting for HTE entirely reverses the empirical findings, while theoretically possible, are rare. There are, however, cases where the magnitude of the coefficient changes substantially; the ratios of FEct estimate to TWFE estimate range between 0.25 to 2.081 in our sample. Notably, both the mean and median of these ratios are close to one. This finding suggests that, although there are noticeable differences in individual cases, TWFE does not systematically under- or over-estimates the ATT. Figure 6(a) also shows that the presence of always-treated units is not the primary driver of these differences. When these units are excluded, the TWFE estimates align closely with the reported estimates.
Figure 6: TWFE versus Imputation Method: Estimates and SEs
(a) Changes in Estimates
(b) Changes in SEs
Note: In subfigure (a), solid circles represent the ratios of the estimates from the imputation method (FEct) to TWFE coefficients with always-treated units removed; hollow circles represent the ratios of reported TWFE coefficients to TWFE coefficients with always-treated units removed. Subfigure (b) shows how changes in SEs lead to changes in $z$-scores for the imputation method: the numerators on both axes are the absolute values of the TWFE coefficients (with always-treated observations removed); the denominators are cluster-bootstrapped SEs from TWFE and from FEct, respectively. Statistically insignificant FEct estimates at the 5% level are painted in red in both subfigures.
Alarmingly, when switching from TWFE to FEct, the number of studies that are statistically insignificant at the 5% level rises from one to 11, constituting 28% of all articles. This loss of significance is, at least in part, driven by larger uncertainty estimates. Interestingly, three out of the ten cases that become insignificant have larger FEct estimates compared to their TWFE counterparts. In Figure 6(b), we plot the ratio of the absolute value of the TWFE estimate to the bootstrapped TWFE SEs against the ratio of the absolute value of the TWFE estimate to the bootstrapped FEct SEs. Thus, this visualization exclusively contrasts the SEs derived from each estimator. In the majority of instances, the SEs for FEct are larger than those from TWFE, and the data points for such cases fall below the 45-degree line. While the gap is generally not substantial, at its maximum, the FEct SE can be more than triple the size of the TWFE SEs. In seven cases, when retaining the TWFE point estimate, adopting the SE from FEct would be sufficient to render the result insignificant. It is not surprising that the loss of statistical significance often occurs in studies where the original results have limited power. One takeaway from this finding is that although it is impossible to definitively tell whether the results from any given model are incorrect (especially without knowing the validity of the PTA), the combination of potential biases from HTE and overconfidence of uncertainty estimates may have produced a significant number of false positives, especially when the original estimates carry a high level of uncertainty.
HTE-robust estimators tend to agree with each other.
[…]

### v3

@@ 5.  Systematic Assessment
The encouraging news is that in over 40% of the studies, we have fairly substantial evidence to attest that either the PTA is plausible or its violation is mild. In these instances, the DTE estimates in the pretreatment periods hover around zero, and formal statistical tests suggest that the remaining imbalance likely results from randomness in data (i.e., the $F$ test does not reject, and the equivalence test does reject using the $|\widehat{ATT}|$ as the equivalence threshold, both at the 5% level). However, for the majority of the remaining studies, except for a few cases where the PTA is highly implausible, the results are less conclusive. We either lack the statistical power or a sufficient number of pre-treatment periods to reliably assess the plausibility of the PTA.
What is much more concerning is that when potential PTA violations are considered, even very mild ones benchmarked against estimated pretrends, most studies lose statistical power. As previously noted, 11 studies lost statistical significance at the 5% level once we apply the imputation estimator. Relaxing the PTA poses an even greater threat. Figure 5(a) shows that the null of no effect is rejected at the 5% level in only in a small minority of the studies—seven among 31 studies (23%) for which such an analysis is feasible—when using the partial identification procedure recommended by Rambachan and Roth (2023) with $\bar{M}=0.5$.
Figure 5: Allowing PTA Violations with Robust Confidence Sets
(a) Robust infernece under PTA violations
(b) Distribution of $\tilde{M}$
Note: Subfigure (a) shows the $p$-values of partially identified ATT estimates using FEct under restricted RM PTA violations with $\bar{M}=0.5$ versus FEct $p$-values assuming the PTA is satisfied ($\bar{M}=0$). Squared root scales are used to facilitate visualization. The red triangles represent studies that lose statistical significance when FEct is used under the PTA. Subfigure (b) is a histogram of $\tilde{M}$, breakdown values of $\bar{M}$ for the 31 studies with more than three pretreatment periods; the pink bar represents 11 studies who ATT estimates are statistically insignificant at the 5% level using FEct.
Figure 5(b) displays the distribution of breakdown values $\tilde{M}$. Among all 31 studies with at least three pretreatment periods, the median is 0.18 and the mean is 0.37. Focusing only on the 20 studies that remain statistically significant at the 5% level with FEct, the median and mean are as low as 0.33 and 0.58, respectively. Although using HTE-robust estimators with robust CSs does not necessarily “overturn” results in most of the paper we reanalyzed, it is evident that the vast majority of studies lack the power to discern the ATT from zero under very mild violations of the key identifying assumption of realistic magnitudes informed by pretrends.
Other issues.
Our reanalysis highlights several additional issues. First, the presence of missing values is widespread. Although most methodological work presumes balanced panels without missing data, in reality, many empirical studies encounter varying degrees of data missingness. We plot patterns of treatment status for each study in the SM. Based on these plots, we see that in some studies the pattern of missingness is either nonrandom or extremely prevalent, which calls into question the reliability and validity of the respective empirical findings.
Second, we perform carryover effect tests for all studies that feature treatment reversals. If this test fails, it suggests that the treatment effects persist beyond the treatment periods. Among 22 articles, five reject the null hypothesis of no carryover effects at 5%. Part of the concern is that imputation methods and DID extensions will use control observations from previously treated units to fit the potential outcome model or as comparisons for treated observations, and if there are carryover effects, then the comparisons will become tainted. LWX (2022) note that the presence of carryover effects for a limited number of periods is less concerning, as researchers can recode treatment to persist for some time after a unit transitions out of treatment. We observe that carryover effects do not substantially alter the findings in most studies. Specifically, when removing two post-treatment periods for studies that reject the null of no carryover effects, the estimates remain similar in magnitude. Additionally, results initially deemed statistically significant continue to maintain their statistical significance (Figure A8 in the SM). Nevertheless, we recommend that researchers make it a practice to check for potential carryover effects, considering the low cost of conducting such tests and adjustments.
Finally, many studies also exhibit sensitivity to model specifications. When we incorporate ULTs into the authors’ original specifications, 41% of studies lose their statistical significance. [^19] Plümper and Troeger (2019) raise a cautionary note against the blanket use of fixed effects models in panel data analysis. They show that if both the treatment and the outcome variables are highly autoregressive, these models may lead to spurious correlations. While the additional parameters reduce efficiency and may lead to biases due to impermissible comparisons (Goodman-Bacon, 2021), these results underscore that a significant number of empirical findings in the literature rely heavily on modeling assumptions. Hence, we recommend researchers carefully assess the robustness of their findings using different model specifications in combination with HTE-robust estimators. We reserve a more careful and rigorous analysis of this issue for future studies.
Relatedly, some studies that we exclude from our sample employ one-way FE or FE at a level different from that at which treatment is assigned. Many of these findings do not hold when we reanalyze them using a TWFE model. We should clarify that this does not imply that the original results are not credible; rather, these models operate under distinct identification assumptions, and there is substantial variation in how much consideration authors give to this point. We notice that some studies do not provide a rationale for their choice to use one-way fixed effects, while others explicitly outline the type of unobserved confounders they intend to control for. In one instance, the authors inaccurately label their specification as a DID design. We emphasize that TWFE and DID are generally not equivalent. This difference becomes even more pronounced when the fixed effects are not assigned at the level of treatment.
Summary.
In Table A1 below, we summarize the main findings of our reanalysis. The numbers represent the proportion of studies in a given journal or across all journals that fall under the respective category. The first two columns relate to the ratio of the FEct estimator to the TWFE estimator, which proxies the consequences of HTE. Across all journals, this ratio is less than 0.8 in 21% of studies and less than 0.5 in 5% of studies, indicating that while HTE might be a significant concern theoretically, its empirical impact may not be as substantial as the existing literature might suggest.
Table 4: Summary of Findings
Consequence of HTE Potential PTA violations Insufficient power
(1) (2) (3) (4) (5) (6) (7) (8)
FEct-TWFE FEct-TWFE $F$ test Placebo test Equiv. test TWFE not FEct not FEct not reject
Journal Ratio $<0.8$ Ratio $<0.5$ reject null reject null not reject null reject null reject null null w/. $\bar{M}=0.5$
APSR (6) 0.00 0.00 0.33 0.33 0.33 0.00 0.17 0.83
AJPS (12) 0.15 0.00 0.17 0.25 0.42 0.00 0.15 0.55
JOP (19) 0.33 0.11 0.33 0.12 0.53 0.06 0.44 0.86
All (37) 0.22 0.05 0.27 0.21 0.45 0.03 0.30 0.74
Note: A null is deemed being rejected if $p<0.05$. In columns (1)–(2), “ratio” refers to the ratio of the FEct estimate to TWFE estimate for the average effect. Four studies with only a single pretreatment period are not included in the summary statistics in columns (3)–(5). Six studies are not included in column (8) because too few pretreatment periods.
Columns (3)–(5) indicate the proportion of studies that have potential PTA violations. The numbers represent the proportion of studies where the null hypothesis is rejected at the 5% significance level for the $F$ test or the placebo test, or not rejected for the equivalence test. Across all journals, around a quarter of studies reject the null hypothesis using the $F$ test, with slightly fewer rejecting when using the placebo test, indicating potential PTA violations. Note, though, that failure to reject may result from insufficient power. In 45% of the cases, the equivalence test fails to reject the null hypothesis using $|\widehat{ATT}|$ as the threshold, implying that for these studies we cannot confidently state that the pretreatment residual averages lie within a narrow range defined by the estimated ATT.
Columns (6) and (7) display the proportion of studies in which the null hypothesis of no effect is not rejected by TWFE and FEct, respectively. When employing the TWFE method, this happens in only one study (3%); however, when the FEct estimator is utilized, this number increases to 30%, suggesting many studies in our sample are potentially underpowered. Moreover, column (8) shows that in 74% of the studies, the robust CSs for the ATT include zero with $\bar{M}=0.5$. This further underscores the fragility of results in the existing literature when we allow very mild violations of the PTA. It is worth noting that in the nine studies with a breakdown value of $\tilde{M}>0.5$, we also observe rejection in the equivalence tests, suggesting a convergence of evidence from multiple statistical tests.
6.  What To Do and (Not) To Do with Causal Panel Analysis
We conclude by sharing practical advice based on both our findings and our experiences conducting a vast number of replications and reanalyses using observational panel data. Table 5 summarizes these lessons. The first consideration is the research design. While fixed effects models allow us to control for certain unobserved confounders, this comes at the cost of strong assumptions about functional form as well as the absence of dynamics between past outcome/covariates and the treatment. If a glaring violation is already known to exist at this stage, a different identification strategy is needed (e.g., Blackwell and Glynn, 2018).
If time-invariant unobserved confounding is a main concern and the PTA or strict exogeneity is plausible, we encourage researchers to plot the data at hand to get a better understanding in the patterns of the treatment status, outcome and missingness. Ideally, treatment status should vary both by unit and time. If the majority of variation occurs over time (across unit) with little or no variation between units at any given time period (or across time within a given unit), the TWFE estimand will be likely dominated by impermissible comparisons and thus potentially highly biased. Furthermore, HTE-robust estimators will estimate the treatment effect using very little data and thus be underpowered. Equally important is the need for researchers to understand the degree and possible origins of data missingness prior to initiating statistical analysis. If missingness does not seem to be random, or if it is too prevalent, leaving insufficient meaningful variation in the data, researchers should consider halting the analysis at this stage. Indeed, proceeding under these circumstances can lead to biased estimates and unreliable conclusions. Just as in the cross-sectional case, plotting the raw data can also help researchers to spot outliers and highly skewed distributions, which may require additional preprocessing.
Table 5: Do’s and Don’ts with Causal Panel Analysis
Do’s Don’ts
Think through research design Start empirical analysis with a research design; clearly specify causal estimands and their corresponding identifying assumptions Start empirical analysis by blindly running regressions; equate designs with outcome models
Estimation Choose HTE-robust estimators and always make the event-study plot Choose models arbitrarily; report regression coefficients only without visualization or diagnostics; equate absence of evidence as evidence of absence.
Diagnostics Conduct both visual and statistical tests, including robust CSs, to gauge whether the PTA and the no-carryover-effect assumption are plausible and the consequences of their potential violations
At the stage of estimation, we recommend choosing an HTE-robust estimator that is suitable for your research context. Although our study reveals that most substantive findings based on TWFE in the existing literature remain unchanged when HTE-robust estimators are employed, this is an empirical observation and not a theoretical guarantee. HTE-robust estimators support different conclusions than the TWFE estimator in a non-negligible minority of our sample, and it is not sufficient to simply hope that one’s own study does not fall within this group. There is a loss of efficiency with these estimators, but we argue this is an acceptable price. Moreover, it is always possible to include a potentially more precise TWFE estimate in addition to the main, HTE-robust estimate when power is a concern and effects are close to homogeneous or heterogeneous effects can be modelled parametrically or semi-parametrically. Diagnostics is crucial. We recommend using the chosen estimator to make the event-study plot, then conducting both a visual inspection and statistical tests, including a sensitivity analysis with robust CSs with different values of $\bar{M}$, to assess the validity of the PTA. Importantly, researchers should refrain from equating the lack of statistical significance for coefficients in the pretrend as evidence for the validity of the PTA.
Finally, if treatment effect heterogeneity is a primary concern in estimating average causal effects, it is sensible to explore the sources of such heterogeneity by summarizing treatment effects across time, unit (or cohort), or pretreatment covariates. It would also be beneficial to model this heterogeneity parametrically or semiparametrically to further improve efficiency.
Appendix A Online Supplementary Materials
  1. A.1.
    Survey of HTE-Robust Estimators
    • A.1.1.
      DID Extensions for the Staggered Setting
    • A.1.2.
      DID Extensions for the General Setting
    • A.1.3.
      Imputation Methods for the General Setting
    • A.1.4.
      PTA and Strict Exogeneity
    • A.1.5.
      Assumptions for Each Estimator
  2. A.2.
    Proof of Remark A.1
  3. A.3.
    Sample and Replicability
    • A.3.1.
      Sample Selection Criteria
    • A.3.2.
      Replicability
  4. A.4.
    Implementation Details
    • A.4.1.
      Summary of Context and Visualization
    • A.4.2.
      Point Estimates
    • A.4.3.
      Dynamic Treatment Effects & Event Study Plots
    • A.4.4.
      Diagnostic Tests
    • A.4.4.
      Robust Confidence Set & Sensitivity Analysis
  5. A.5.
    More Replication Results
    • A.5.1.
      Reported vs TWFE and FEct Estimates
    • A.5.2.
      Inferential Methods
    • A.5.3.
      Alternative Specifications
    • A.5.4.
      Placebo Tests
    • A.5.5.
      Carryover Effects
    • A.5.6.
      Summary of Findings
A.1 Survey of HTE-Robust Estimators
Scholars have proposed a number of novel estimators that relax TWFE assumptions and allow for HTE. We discuss several of them below. Broadly, we can categorize these estimators along two dimensions: (1) estimation strategy and (2) applicable settings. Along (1), we divide estimators into two groups. We call one group of methods “DID extensions,” which use local, $2\times 2$ DIDs between treated and control observations as building blocks, and the other “imputation methods,” which impute counterfactual outcomes using an explicit outcome model (in particular, the TWFE model) that is fit globally on all available control observations. We see the former as direct extensions to DID, while the latter embed DID’s functional form assumptions in their outcome models. For (2), estimators either are suited only to the staggered setting (which includes the classic DID setting) where treatment is an absorbing state or can accommodate treatment reversals. Those suited to the latter are also suited to the former, which is just a special case of the latter. The reverse is not true.
In the following subsections, which are organized by this typology, we introduce and compare several recently introduced HTE-robust estimators. Although these estimators all relax the TWFE assumption of homogeneous effects, they do not absolve us of needing the parallel trends assumption (PTA) or strict exogeneity. These estimators can, however, estimate dynamic treatment effects, which in turn allow us to assess the validity of parallel trends by testing for pretrends.
A.1.1 DID Extensions for the Staggered Setting
We first introduce a set of estimators, each constructed from local $2\times 2$ DID estimates, that are suitable only for the staggered adoption setting. The general strategy of these estimators is to estimate the dynamic cohort average treatment effect on the treated (CATT), $\delta_{g,l}$, for each cohort $g$ and for each period since treatment adoption $l$ using a valid $2\times 2$ DID. By valid, we mean that the DID consists of (1) a pre-period and a post-period and (2) a treated group and a comparison group. The pre-period is such that all observations in both groups are in control, and the post-period is such that observations from the treated group are in treatment and the observations from the comparison group are in control. The choice of comparison group is what primarily distinguishes estimators in this category. To obtain higher-level averages, we then average over our estimates of $\delta_{g,l}$ using appropriate, non-negative weights.
Sun and Abraham (2021) propose an interaction-weighted (iw) estimator that is a weighted average of CATT estimates obtained from a TWFE regression with cohort dummies fully interacted with indicators of relative time to the treatment’s onset. Specifically,
$Y_{i,t}=\alpha_{i}+\lambda_{t}+\sum_{g\notin\mathcal{C}}\sum_{l\neq 0}\delta_{g,l}\mathbf{1}\{E_{i}=g\}\cdot\mathbf{1}\left\{K_{i,t}=l\right\}+\epsilon_{i,t},\vspace{-0.5em}$ (A1)
where $\mathcal{C}$ is some set of reference cohorts and $K_{i,t}$ is similarly defined as in the main text. Equivalently, each estimate of $\delta_{g,l}$ from equation A1 can can be characterized as a difference in the average changes in outcome from some fixed pre-period $s<g$ to $l$ periods since $g$ between the treated cohort $g$ and comparison cohorts in $\mathcal{C}$:
$\hat{\delta}(g,l)=\frac{1}{|\{i:E_{i,t}=g\}|}\sum_{i:E_{i,t}=g}(Y_{i,g+l}-Y_{i,s})-\frac{1}{|\{i:E_{i,t}\in\mathcal{C}\}|}\sum_{i:E_{i,t}\in\mathcal{C}}(Y_{i,g+l}-Y_{i,s}),$
The authors recommend using $\mathcal{C}=\{\sup_{i}E_{i,t}\}$, which is either the never-treated cohort or (if none exists) the last-treated cohort. The estimator then weights $\hat{\delta}_{g,l}$ by the sample share of each cohort $\hat{w}_{g}$ before taking some average thereof. For example, the dynamic treatment effects (DTE) from relative period $l$ between $-a$ and $b$ can be estimated from
$\hat{\delta}^{IW}_{l}=\sum_{g}\hat{w}_{g}\hat{\delta}_{g,l},\quad a\leq l\leq b,\vspace{-0.5em}$
and the ATT up to $b$ periods after the treatment’s onset from
$\hat{\delta}^{IW}=\frac{1}{b}\sum_{1\leq l\leq b}\sum_{g}\hat{w}_{g}\hat{\delta}_{g,l}.$
The authors note that their estimator can be extended to include covariates, but also that this may require additional functional form assumptions.
Using the same general strategy, Callaway and Sant’Anna (2021) propose doubly robust estimators that directly incorporate pretreatment covariates. These estimators, which we collectively refer to as csdid, use either never-treated ($\hat{\delta}^{CS-dr}_{nev}$) or not-yet-treated units ($\hat{\delta}^{CS-dr}_{ny}$) as the comparison group. $\hat{\delta}^{CS-dr}_{nev}$ uses the same comparison group as iw when a never-treated cohort exists, whereas $\hat{\delta}^{CS-dr}_{ny}$ differs and uses all untreated observations of later adopters (including the never-treated) as potential controls for early adopters. Besides the choice of comparison cohort, these estimators both differ from the iw estimator in that they allow the user to condition on pretreatment covariates using both an explicit outcome model and inverse propensity score weighting (IPW).A1A1A1The IPW estimator proposed by Strezhnev (2018) is similar to $\hat{\delta}^{CS-dr}_{ny}$. One small difference is that $\hat{\delta}^{CS-dr}_{ny}$ allows more complex outcome modeling than a simple before-and-after estimator. If either the outcome model or the propensity score model is correct, the estimators will be consistent.
A.1.2 DID Extensions for the General Setting
The next group of estimators we discuss also use local DIDs as building blocks, but estimators in this group can accommodate treatment reversals. The general strategy is once again to use valid $2\times 2$ DIDs, but this time the goal is to estimate the DTE $\delta_{l}$ for all treated units some number of periods since treatment $l$—cohorts are no longer defined, since treatment reversals make it insensible to group units by their time of treatment adoption. The literature has effectively proposed one common strategy of selecting a comparison group, which is to match treated and control observations belonging to units with the same treatment history.
IKW (2023) propose one such estimator. Formally, to estimate the ATT, we first define a matched set for each observation $(i,t)$ satisfying $D_{i,t}=1$ and $D_{i,t-1}=0$,
$\mathcal{M}_{i,t}=\bigg{\{}i^{\prime}:i^{\prime}\neq i,D_{i^{\prime},t}=0,D_{i^{\prime},t^{\prime}}=D_{i,t^{\prime}}\;\forall t^{\prime}\in\{t-1,t-2,\dots,t-a\}\bigg{\}},\vspace{-1em}$
where $a$ is the number of periods on which we wish to match treatment histories. The authors also propose “refining” the matched set to incorporate other pretreatment covariates and past outcomes. We do not further discuss refinement for a more seamless comparison with other estimators and refer interested readers to the original paper. Without refinement and fixing the number of periods $a$ on which to match, the proposed estimator for the DTE $l$ periods since treatment $\delta_{l}$ is,
$\hat{\delta}^{PM}_{l,a}=\frac{\sum_{t=a}^{T-l+1}\sum_{i=1}^{N}G_{i,t}\hat{\delta}_{l}^{(i,t)}}{\sum_{t=a+1}^{T-l}\sum_{i=1}^{N}G_{i,t}},$
where $G_{i,t}=\mathbf{1}\left\{|\mathcal{M}_{i,t}|>0\right\}D_{i,t}(1-D_{i,t-1})$ is equal to 1 if and only if the observation $(i,t)$ switches into treatment at time $t$ and has a non-empty matched set (and is zero otherwise) and $\hat{\delta}_{l}^{(i,t)}=(Y_{i,t-1+l}-Y_{i,t-1})-\sum_{i^{\prime}\in\mathcal{M}_{i,t}}\frac{1}{|\mathcal{M}_{i,t}|}(Y_{i^{\prime},t-1+l}-Y_{i^{\prime},t-1})$ is the local DID obtained from the pre- and post-periods $t-1$ and $t-1+l$, respectively, the treatment “group” consisting of just $(i,t)$, and the comparison group consisting of the matched set for $(i,t)$. To then obtain an estimate for the DTE $\hat{\delta}_{l}$, we then average over all $\hat{\delta}_{l}^{(i,t)}$ such that $(i,t)$ Essentially, the strategy is to average over the estimates of the DTE for all units that switch into treatment at $t$ (if there are any) for each time period $t$, and then to average across all time periods for which we can obtain an estimate.A2A2A2Note that, without refinement, all treated observations with the same treatment history share the same matched set, so we can group these observations together and rewrite the inner sum to instead be over all possible treatment histories. We can thus also express the inner sum of the numerator as a weighted sum of local DIDs using a slightly different treatment group—all treated observations with the treatment history—where the weights are proportional to the size of said group. If the goal is to estimate the average effect of treatment reversal (ART), we then analogously defined matched sets for each observation $(i,t)$ satisfying $D_{i,t}=1$ and $D_{i,t-1}=0$, $\mathcal{M}_{i,t}=\{i^{\prime}:i^{\prime}\neq i,D_{i^{\prime},t}=1,D_{i^{\prime},t^{\prime}}=D_{i,t^{\prime}}\;\forall t^{\prime}\in\{t-1,t-2,\dots,t-a\}\}$. We use $\hat{\delta}^{PM-ART}_{l,a}$ to denote the resulting estimator.
Interestingly, several DID extensions can be viewed as special cases of PanelMatch.
Remark A.1 (Relation between $\hat{\delta}_{1,1}^{PM}$ without refinement and $\hat{\delta}^{M}$).
Assume we have a balanced panel of units, i.e. every unit $i$ is observed at every time period $t$. For the special case when we match on only one period ($a=1$) and are estimating the contemporaneous treatment effect ($l=1$), without refinement, a weighted average of the PanelMatch estimators for the ATT and ART is equivalent to the multiple DID estimator proposed by de Chaisemartin and D’Haultfœuille (2020), or $\hat{\delta}^{M}$, when there exists a ‘stable’ group (i.e., whenever there is a unit switching into or out of treatment, there is at least one other unit staying in control or treatment; see the next section for a formal statement of this assumption), where the weights are the proportion of ‘switchers’ that are “joiners” versus “leavers.” That is, if we do not refine the matched set, then $\frac{N_{J}}{N_{S}}\hat{\delta}^{PM}_{1,1}+\frac{N_{L}}{N_{S}}\hat{\delta}^{PM-ART}_{1,1}=\hat{\delta}^{M}$, where $N_{J},N_{L},$ and $N_{S}$ are the numbers of joiners, leavers, and switchers. The proof is in the next section. This observation allows us to appeal to the results that de Chaisemartin and D’Haultfœuille (2020) prove about $\hat{\delta}^{M}$. Minor adjustments of their proofs will give us that, under some typical assumptions (the details of which we provide later in this section), $\hat{\delta}_{1,1}^{PM}$ without refinement is asymptotically normal, unbiased, consistent for the average contemporaneous treatment (reversal) effect on the treated.
Remark A.2 (Equivalence of PanelMatch and csdid without covariate adjustment).
Again assume we have a balanced panel of units. If we use a simple difference in means as the outcome model for csdid and employ uniform propensity score weights (i.e., do not adjust for covariates), then csdid is equivalent to PanelMatch with an arbitrary number of lags and without refinement (in the staggered setting). This follows from the facts that in the staggered setting, for any time period $t$: (1) Any observation belonging to a unit that switches into treatment at time $t$ (‘switchers’) must have been under control for the periods $11,\dots t-$ 1; and (2) all control observations must belong to units that have been under control for the periods $11,\dots t-$ 1 (i.e., they have the same treatment history as switchers). Thus, the matched set will always include all units under control (all “not-yet treated” units).
A.1.3 Imputation Methods for the General Setting
The last class of estimators we discuss no longer directly take the difference between differences; instead, they take the difference of the observed outcome and an imputed counterfactual outcome (for treated observations)—the before-and-after difference is embedded in the functional form assumption used to impute treated counterfactuals. Under strict exogeneity or a stronger version of the PTA, the imputation method allows researchers to make inferences about the ITE of treated observations, $\tau_{i,t},\forall(i,t)\;s.t.\;D_{i,t}=1$, the most fine-grained estimand (e.g., Bai and Ng, 2021b).
BJS (2023) propose an “imputation procedure” that first imputes the counterfactual outcomes for treated units based on the outcome model,
$Y_{i,t}=A_{i,t}^{\prime}\lambda_{i}+X_{i,t}^{\prime}\beta+D_{i,t}\Gamma_{i,t}^{\prime}\theta+\epsilon_{i,t},\vspace{-1em}$
and then estimates the treatment effect for treated observations with the difference between their observed and their imputed counterfactual outcomes. That is, first use only the untreated observations $\{(i,t):D_{i,t}=0\}$ to estimate $\lambda_{i}$ and $\beta$ (by $\hat{\lambda}_{i}$ and $\hat{\beta}$) using OLS on the regression $Y_{i,t}=A_{i,t}^{\prime}\lambda_{i}+X_{i,t}^{\prime}\beta+\varepsilon_{i,t}$. Then, for each treated observation, set $\hat{Y}_{i,t}^{BJS}(0)=A_{i,t}^{\prime}\hat{\lambda}_{i}+X_{i,t}^{\prime}\hat{\beta}$ and estimate the ITE as $\hat{\delta}^{BJS}_{i,t}=Y_{i,t}-\hat{Y}_{i,t}^{BJS}(0)$. We can then combine these ITE estimates to estimate aggregate quantities, including the ATT and dynamic effects.
LWX (2022) refer to imputation-based estimators as “counterfactual estimators” and discuss several such estimators. LWX (2022) consider a class of outcome models of the form $Y_{i,t}(0)=f(X_{i,t})+h(U_{i,t})+\varepsilon_{i,t}$, where $f(\cdot)$ and $h(\cdot)$ are known parametric functions, $X_{i,t}$ is observed, and $U_{i,t}$ is unobserved (whereas in BJS (2023), both $X_{it}$ and $A_{it}$ are observed). Note that this framework subsumes the TWFE outcome model as we can model $Y_{i,t}(0)=X_{i,t}^{\prime}\beta+\alpha_{i}+\xi_{t}+\varepsilon_{i,t}$. We can then use an estimation procedure similar to the one in BJS (2023). LWX (2022) call this estimator the fixed effect counterfactual (FEct) estimator, $\hat{\delta}^{fect}$ (for the ATT) or $\hat{\delta}^{fect}_{l}$ (for the dynamic effects).
A.1.4 PTA and Strict Exogeneity
We first discuss and compare the key identification assumptions required by each method.
Sun and Abraham (2021) define potential outcomes based on treatment history and assume parallel trends for the never-treated potential outcome $Y_{i,t}(\infty)$ of the comparison group: $\mathbb{E}[Y_{i,t}(\infty)-Y_{i,s}(\infty)|E_{i,t}=e]$ is the same for all $s\neq t$ and $e\in\mathop{\rm supp}E_{i,t}$.A3A3A3We state the unconditional versions of these assumptions for simplicity. Call this assumption “parallel trends A.” Callaway and Sant’Anna (2021) similarly assume parallel trends for the comparison group, but define potential outcomes based on current treatment status. As a result, the statement of the assumption becomes, for all $g$, $\mathbb{E}[Y_{i,t}(0)-Y_{i,t-1}(0)|E_{i}=g]=\mathbb{E}[Y_{i,t}(0)-Y_{i,t-1}(0)|E_{i,t}\in\mathcal{C}]$ for each $t\geq\max\{2,g\}$. Call this version “parallel trends B.”
de Chaisemartin and D’Haultfœuille (2020) assume both “strong exogeneity” and “common trends.” They define the former as, $\mathbb{E}[Y_{i,t}(d)-Y_{i,t-1}(d)|\{D_{i,t}\}_{t=1}^{T}]=\mathbb{E}[Y_{i,t}(d)-Y_{i,t-1}(d)]$ for all $i$, all $t\geq 2$, and all $d\in\{0,1\}$.A4A4A4The actual assumption is that this equality holds for all groups $g$, where $g$ is the level of the fixed effects, which may be at a higher level than the unit level (e.g., if $i$ indexes cities, $g$ might be counties or states/provinces). For consistency and simplicity, we assume that this is equal to the unit level in our discussion (i.e., unit fixed effects). The common trends assumption requires that this last quantity — that is, $\mathbb{E}[Y_{i,t}(d)-Y_{i,t-1}(d)]$ — does not vary across $i$ for all $t\geq 2$ and $d\in\{0,1\}$. Combining these two assumptions, we can instead write that $\mathbb{E}[Y_{i,t}(d)-Y_{i,t-1}(d)|\{D_{i,t}\}_{t=1}^{T}]=\mathbb{E}[Y_{j,t}(d)-Y_{j,t-1}(d)]$ for all $j$ (including $j=i$), all $i$, all $t\geq 2$, and all $d\in\{0,1\}$. Call this combined version of the assumptions “parallel trends C.” Like Sun and Abraham (2021), IKW (2023) define potential outcomes in terms of treatment histories. IKW (2023) do not, however, assume staggered adoption, and so a much wider range of treatment histories are possible. The comparison group is also substantially different. The latter compares units that switch into treatment with those that stay in control and asks that their respective trends be parallel: $\mathbb{E}[Y_{i,t+l}(D_{i,t}=0,D_{i,t-1}=0,\{D_{i,t-s}\}_{s=2}^{a})|D_{i,t}=1,D_{i,t-1}=0]=\mathbb{E}[Y_{i,t+l}(D_{i,t}=0,D_{i,t-1}=0,\{D_{i,t-s}\}_{s=2}^{a})|D_{i,t}=0,D_{i,t-1}=0]$. Call this assumption “parallel trends D.”
Recall that the imputation estimators connect to DID in a less direct way, which in turn implies different assumptions: They assume a TWFE model for untreated potential outcomes, which requires mean independence for all pairs of units $i,j$ and all pairs of time periods $t,s$. For example, BJS (2023) define a version of parallel trends as $\mathbb{E}[Y_{i,t}(0)-Y_{i,s}(0)]=\mathbb{E}[Y_{j,t}(0)-Y_{j,s}(0)]$ for all $i,j$ and all $t,s$. The estimator from BJS (2023) does not require this to hold, instead requiring a weaker assumption, $\mathbb{E}[Y_{i,t}(0)]=A_{i,t}^{\prime}\lambda_{i}+X_{i,t}^{\prime}\delta$ for all $i,t$. Note that this assumption implies that each idiosyncratic error is zero in expectation, and thus we refer to this assumption as “outcome model and mean-zero errors.” $\hat{\delta}^{fect}$ from LWX (2022) requires strict exogeneity and the TWFE outcome model, which together imply the PTA defined by BJS (2023).
A.1.5 Assumptions for Each Estimator
Next, we provide a fuller account of all assumptions invoked by each method.
Sun and Abraham (2021)
  • Parallel trends A; and
[…]